# A Practical Investigation of Spatially-Controlled Image Generation with Transformers

### v1

@@ Abstract
Enabling image generation models to be spatially controlled is an important area of research, empowering users to better generate images according to their own fine-grained specifications via e.g. edge maps, poses. Although this task has seen impressive improvements in recent times, a focus on rapidly producing stronger models has come at the cost of detailed and fair scientific comparison. Differing training data, model architectures and generation paradigms make it difficult to disentangle the factors contributing to performance. Meanwhile, the motivations and nuances of certain approaches become lost in the literature. In this work, we aim to provide clear takeaways across generation paradigms for practitioners wishing to develop transformer-based systems for spatially-controlled generation, clarifying the literature and addressing knowledge gaps. We perform controlled experiments on ImageNet across diffusion-based/flow-based and autoregressive (AR) models. First, we establish control token prefilling as a simple, general and performant baseline approach for transformers. We then investigate previously underexplored sampling time enhancements, showing that extending classifier-free guidance to control, as well as softmax truncation, have a strong impact on control-generation consistency. Finally, we re-clarify the motivation of adapter-based approaches, demonstrating that they mitigate “forgetting” and maintain generation quality when trained on limited downstream data, but underperform full training in terms of generation-control consistency. Code will be released upon publication.
@@ 2 Preliminaries
results in simple linear combinations of the score estimates $\bm{s}_{\bm{\theta}}(\bm{z}_{t},t)\approx\nabla_{\bm{z}_{t}}\log\;p(\bm{z}_{t})$ (see Eq. 4) and pre-softmax logits $\bm{v}$. Note that the normalising constant $Z$ is naturally dealt with for diffusion due to the score being a gradient, whilst in VAR it is taken care of by the softmax denominator. Note that setting $\gamma_{y}=0$ recovers the unconditional model $p_{\bm{\theta}}(\bm{z})$, whilst $\gamma_{y}=1$ recovers the conditional one $p_{\bm{\theta}}(\bm{z}|\bm{y})$. $\gamma_{y}$ is typically set to be $\geq 1$ in order to improve generation quality.
@@ 3 Evaluation Setup
In order to perform a clean and fair comparison and avoid potential confounding factors related to data quality (during both pre-training and control finetuning), we train and evaluate on class-conditioned ImageNet (Deng et al., 2009), a well-established benchmark for image generation. ImageNet is at a scale ($\sim$1.2M samples, 256$\times$256 resolution, baseline models with $10^{8}\sim 10^{9}$ parameters) that is close to deployed real-world image generation models, whilst still having manageable training and evaluation costs for an academic compute budget. We train and evaluate at a resolution of 256$\times$256 over two distinct spatial control conditions, which we extract from the ImageNet dataset on the fly during training and evaluation: 1) Canny edge maps (Canny, 1986), using kornia with fixed thresholds $(0.1,0.2)$, and 2) dense depth maps using DPT-Hybrid (Ranftl et al., 2021).
@@
For most evaluations we generate 10K samples for evaluation, conditioned on controls extracted from the first 10 images of each of the 1000 classes in the ImageNet validation dataset. In a few cases, to compare with the literature, we generate using controls from all 50K validation images. We use fixed random seeds. We measure image generation performance using three metrics. Fréchet Inception Distance (FID) (Salimans et al., 2016) measures the distributional distance between the data distribution and generated samples, whilst Inception Score (IS) (Heusel et al., 2017) measures the diversity and clarity of generations. These are widely used image quality metrics. Generation-control consistency (Li et al., 2024a) intuitively measures how closely generations adhere to spatial control, via similarity between the original control and the control extracted from the generated image. Following Li et al. (2024a), for Canny edge maps we compare binary pixels using F1 score $\uparrow$, whilst for depth maps we use root mean square error $\downarrow$ (RMSE) for pixel values in $[0,255]$. Consistency is a naturally desirable property, i.e. to have the generation closely match the provided conditioning, and is especially motivated by applications with tight spatial tolerances such as industrial design.
@@
We compare generation throughput in images per second (img/s) and latency in milliseconds (ms) by directly measuring time in python. Inference is performed on a single NVIDIA Tesla V100-32GB-SXM2. We do not include the decoder to the pixel space, or the extraction and encoding of control data to the latent space.
@@ 4 Prefilling – a Baseline for Transformer-Based Controllable Generation
model #params
@@
VAR-d16 CFG=2.5, temp=1, top-$p$=0.96, top-$k$=900 depth prefill 310M 3.30 207 5.95 165 31.3 0.31 12
@@
SiT-XL/2 CFG=1.5, Euler ODE steps=64 depth prefill 675M 2.78 218 5.43 170 29.3 3.4 0.35
@@
Figure 2: Left: Examples of conditional generations using prefilling. Additional examples can be found in Appendix B. Right: Results for our prefill baseline where each model is finetuned from an ImageNet-pretrained base model. “*” indicates results copied from Li et al. (2025; 2024b). With default sampling parameters, prefilling is able to generate higher quality images than recent transformer-based approaches with comparable control consistency. We additionally note the relatively low overhead for SiT, as KV-caching means the control tokens use up only a single forward pass. Quantising using the VQVAE hurts control consistency for VAR, without much effect on generation quality.
We investigate two recent transformer-based approaches for image generation, to cover both diffusion-based and (comparatively underexplored) autoregressive methods: 1) SiT (Ma et al., 2024), which trains a diffusion transformer (DiT) architecture (Peebles & Xie, 2023) using flow matching, and 2) VAR, which autoregressively generates image representations of increasing spatial scale (Tian et al., 2024). Fig. 1 illustrates our prefill baseline for both cases.
Implementation and experimental details.
To implement the above prefilling approach (Fig. 1), we simply fine-tune the existing architectures with additional conditioning tokens, which the generation tokens attend to via a block-causal mask. We consider VAR-d16 and SiT-XL/2, initialising both models with ImageNet-pretrained checkpoints. [^3] VAR: https://github.com/FoundationVision/VAR, SiT: https://github.com/sihyun-yu/REPA (Yu et al., 2025) The base architectures are left unchanged. For VAR we add new learnable positional embeddings for the control tokens and a shared additive “level” embedding for all control tokens, whilst for SiT we only add a new “level” embedding for the control tokens, reusing the original sinusoidal positional embeddings. For SiT, we set diffusion time $t=0$ (clean data) for control tokens. We finetune for 10 epochs with control conditioning with batch size 256 ($\sim$50K iterations) using the original optimisers and hyperparameters. Following the original papers we linearly increase guidance scale $\gamma_{y}$ from zero over generation scales for VAR, whilst keeping it constant for SiT. CFG is performed by only omitting class conditioning and not control conditioning. VAR and SiT use different image crops, leading to small differences in visualisations.
Results.
@@
Takeaway: Prefilling (Fig. 1) is a simple, general and effective baseline for spatial control with transformers.
We note that our result goes against the results in ControlAR (Li et al., 2025) that suggest that prefilling is a poor choice. We hypothesise that Li et al. (2025)’s choices to 1) vector quantise the control input, destroying information, and 2) enforce a triangular causal attention mask, limiting attention between control tokens, may have hurt performance. We perform our own experiment with VAR’s VQVAE (Fig. 2) confirming the negative effect of quantising the control input. Although avoiding quantisation may sound obvious, AR models are often trained using token indices as inputs by default (Sun et al., 2024), meaning that “turning off” quantisation may lead to additional implementation complexity.
@@ 5.1 Control Guidance (ctrl-G)
Consider the idea of CFG introduced in Eq. 6; it can be easily extended to the spatial control condition $\bm{c}$, to concentrate sampling density in regions where $p(\bm{c}|\bm{z})$ is implicitly higher, potentially improving control consistency,
@@
Applying ctrl-G to SiT quickly runs into an issue – generations become highly saturated, shown in Fig. 3. Thankfully, Sadat et al. (2025) recently propose projected guidance (proj-G) to deal with contrast/saturation issues arising in large CFG scenarios. Applying their approach to ctrl-G, we get,
@@
Takeaway: For flow/diffusion models, control guidance (Eq. 9) leads to oversaturation. Applying projected guidance (Sadat et al., 2025) (i.e. only using the orthogonal component Eq. 11) assuages this issue.
@@ Results.
Fig. 4 shows how image generation quality and control consistency vary with both CFG ($\gamma_{y}$) and ctrl-G ($\gamma_{c}$). Compared to the default settings used in Fig. 2, we find that increasing CFG generally improves both IS and FID (although too high a scale may hurt FID), but slightly decreases consistency. Interestingly, in the case of VAR, higher CFG improves FID over the no-control model ($\sim 5.7$ to $\sim 5.1$), demonstrating that providing a weaker model with spatial cues can boost generation quality. On the other hand, ctrl-G is able to drastically improve consistency, but this is traded off against worse generation quality. This is visualised in the left of Fig. 3, where we see that the canny edge map is better conformed to, but at the cost of over-emphasised edges and incongruous generation for larger ctrl-G. Intuitively, ctrl-G only targets the probability of the control given the generation $p(\bm{c}|\bm{z})$ (Eq. 8), which may not align with image quality. This trade-off aligns with the results reported for image editing in (Brooks et al., 2023). We note that the F1 values of $\sim$50 for SiT are significantly higher than the values commonly found in the literature (Li et al., 2025; Xiao et al., 2025), demonstrating the large effect that this generation-time adjustment has on consistency. Additionally, ctrl-G incurs the extra inference cost of needing to run the no-control model $p_{\bm{\theta}}(\bm{z})$, reducing latency and throughput: [^6] We use the pretrained ImageNet model for $p_{\bm{\theta}}(\bm{z})$ in series with the control-prefill model. Better latency and parameter efficiency could be achieved by training the control-prefill model with empty control $\bm{c}=\varnothing$ and parallelising like CFG, although the FLOPs increase would still remain.
@@
Finally we note that in ControlVAR (Li et al., 2024b), when performing guidance (with class and control conditioning), all guidance scales $\gamma$ are set to be equal. However, Fig. 4 shows that this setting may lead to excessive degradation in generation quality, demonstrating the importance of separately tuning $\gamma_{y}$ (CFG) and $\gamma_{c}$ (ctrl-G).
@@ 5.2 Sampling Distribution Truncation
$\pi^{\text{top-}p}_{k}=\frac{\pi_{k}}{\sum_{i\in V_{p}}\pi_{i}}\text{ if }k\in V_{p}\text{,~{}~{}otherwise }0\quad\quad\text{where }V_{p}=\text{the smallest set of indices }V\text{ s.t. }\sum\nolimits_{i\in V}\pi_{i}\geq p~{},$ (13)
effectively discarding the tail $1-p$ of the softmax. Intuitively, one would expect the most probable vector codes, predicted by the model’s softmax at each token location, to best conform to the spatial control condition.
@@ Results.
Figs. 3, 5 and 8 show how top-$p$ and temperature scaling affect generation quality and control consistency alongside CFG. We see that CFG behaves similarly to Fig. 4, generally improving generation quality and slightly hurting consistency. Notably, truncation via top-$p$ is able to improve both generation quality and control consistency (although excessive truncation seems to hurt FID likely due to reduced diversity). Fig. 8 (Appendix A) further shows that the consistency improvements can be stacked with ctrl-G. This demonstrates that top-$p$ is an important parameter to adjust for softmax-based generative models, providing improvements in all fronts without any additional cost. In contrast, temperature scaling is unable to meaningfully improve consistency, leading to a small increase for canny, and no improvement for depth, whilst simultaneously hurting generation quality. Compared to the results for ctrl-G, this suggests that naively scaling the score (without changing its direction) is insufficient to improve consistency. One possible intuition, is that score temperature scaling occurs uniformly over the spatial dimensions, i.e. it is a global transformation, and so it may have limited spatial/local influence, whilst ctrl-G varies in space. On the other hand, top-$p$ explicitly truncates at each token location, allowing it to influence spatial consistency.
@@
We finally note that Figs. 4, 5 and 8 demonstrate that meaningfully improved performance can be easily obtained by the sampling adjustments explore in this section (compared to the default settings used in Fig. 2).
@@ 6.1 Adapter Architectures
Add adapter. Illustrated in the right of Fig. 6, this approach adapts the ControlNet (Zhang et al., 2023) architecture to transformers, like in Chen et al. (2024a), training a copy of the base model to additively inject spatial conditioning information at each layer, via zero-initialised linear layers, whilst the original weights are performing generation. Inference for each generation token is matched by inference on a corresponding control token.
Attention/prefill adapter. Illustrated in the centre of Fig. 6, this approach retains the prefill paradigm of conditioning from the baseline, and trains a separate copy of the transformer blocks to be used during condition prefilling, leaving generation to the original model weights. Although the adapter and base model have different attention weight matrices, generation tokens are able to attend back to the keys and values produced by the control adapter.

### v2

@@ Abstract
Enabling image generation models to be spatially controlled is an important area of research, empowering users to better generate images according to their own fine-grained specifications via e.g. edge maps, poses. Although this task has seen impressive improvements in recent times, a focus on rapidly producing stronger models has come at the cost of detailed and fair scientific comparison. Differing training data, model architectures and generation paradigms make it difficult to disentangle the factors contributing to performance. Meanwhile, the motivations and nuances of certain approaches become lost in the literature. In this work, we aim to provide clear takeaways across generation paradigms for practitioners wishing to develop transformer-based systems for spatially-controlled generation, clarifying the literature and addressing knowledge gaps. We perform controlled experiments on ImageNet across diffusion-based/flow-based and autoregressive (AR) models. First, we establish control token prefilling as a simple, general and performant baseline approach for transformers. We then investigate previously underexplored sampling time enhancements, showing that extending classifier-free guidance to control, as well as softmax truncation, have a strong impact on control-generation consistency. Finally, we re-clarify the motivation of adapter-based approaches, demonstrating that they mitigate “forgetting” and maintain generation quality when trained on limited downstream data, but underperform full training in terms of generation-control consistency. Code: https://github.com/guoxoug/transformer-imagenet-ctrl.
@@ 2 Preliminaries
results in simple linear combinations of the score estimates $\bm{s}_{\bm{\theta}}(\bm{z}_{t},t)\approx\nabla_{\bm{z}_{t}}\log\;p(\bm{z}_{t})$ (see Eq.˜4) and pre-softmax logits $\bm{v}$. Note that the normalising constant $Z$ is naturally dealt with for diffusion due to the score being a gradient, whilst in VAR it is taken care of by the softmax denominator. Note that setting $\gamma_{y}=0$ recovers the unconditional model $p_{\bm{\theta}}(\bm{z})$, whilst $\gamma_{y}=1$ recovers the conditional one $p_{\bm{\theta}}(\bm{z}|\bm{y})$. $\gamma_{y}$ is typically set to be $\geq 1$ in order to improve generation quality.
@@ 3 Evaluation Setup
In order to perform a clean and fair comparison and avoid potential confounding factors related to data quality (during both pre-training and control finetuning), we train and evaluate on class-conditioned ImageNet (Deng et al., 2009), a well-established benchmark for image generation. ImageNet is at a scale ($\sim$1.2M samples, $\times256$ resolution, baseline models with $10^{8}\sim 10^{9}$ parameters) that is close to deployed real-world image generation models, whilst still having manageable training and evaluation costs for an academic compute budget. We train and evaluate at a resolution of $\times256$ over two distinct spatial control conditions, which we extract from the ImageNet dataset on the fly during training and evaluation: 1) Canny edge maps (Canny, 1986), using kornia with fixed thresholds $(0.1,0.2)$, and 2) dense depth maps using DPT-Hybrid (Ranftl et al., 2021).
@@
For most evaluations we generate 10K samples for evaluation, conditioned on controls extracted from the first 10 images of each of the 1000 classes in the ImageNet validation dataset. In a few cases, to compare with the literature, we generate using controls from all 50K validation images. We use fixed random seeds. We measure image generation performance using three metrics. Fréchet Inception Distance (FID) (Salimans et al., 2016) measures the distributional distance between the data distribution and generated samples, capturing both realism and diversity, whilst Inception Score (IS) (Heusel et al., 2017) measures the (class-)diversity and clarity of generations, reflecting how confidently a pretrained classifier can assign labels and how varied those labels are across samples. These are widely used image quality metrics. Generation-control consistency (Li et al., 2024a) intuitively measures how closely generations adhere to spatial control, via similarity between the original control and the control extracted from the generated image. Following Li et al. (2024a), for Canny edge maps we compare binary pixels using F1 score $\uparrow$, whilst for depth maps we use root mean square error $\downarrow$ (RMSE) for pixel values in $[0,255]$. Consistency is a naturally desirable property, i.e. to have the generation closely match the provided conditioning, and is especially motivated by applications with tight spatial tolerances such as industrial design. Additionally we emphasise the following:
Takeaway: Metrics for generation quality (FID, IS) and control consistency (F1, RMSE) quantify certain aspects of performance; however, practitioners’ requirements subjectively depend on their individual use cases (e.g. prioritise consistency vs aesthetics, don’t care about diversity). Readers should approach the quantitative results presented in this paper considering both the nature of the metrics reported and their own individual requirements.
@@
We compare generation throughput in images per second (img/s) and latency in milliseconds (ms) by directly measuring time in python. Inference is performed on a single NVIDIA Tesla V100-32GB-SXM2. We do not include the decoder to the pixel space, or the extraction and encoding of control data to the latent space. This is to isolate the impact on the transformer component of the model, as that is the design space we are exploring.
Limitations.
Although the aim of the above experimental setup is to provide practically useful and generalisable takeaways, we also qualify that it is not comprehensive. For example, other than the additional experiment in Appendix˜C, we do not perform experiments on text-to-image models. Neither do we directly evaluate raster-scan AR image generation models. Readers should take this into consideration when interpreting our results/contributions for their own use.
@@ 4 Prefilling – a Baseline for Transformer-Based Controllable Generation
We investigate two recent transformer-based approaches for image generation, to cover both diffusion-based and (comparatively underexplored) autoregressive methods: 1) SiT (Ma et al., 2024), which trains a diffusion transformer (DiT) architecture (Peebles & Xie, 2023) using flow matching, and 2) VAR, which autoregressively generates image representations of increasing spatial scale (Tian et al., 2024). Fig.˜1 illustrates our prefill baseline for both cases.
Implementation and experimental details.
To implement the above prefilling approach (Fig.˜1), we simply fine-tune the existing architectures with additional conditioning tokens, which the generation tokens attend to via a block-causal mask. We consider VAR-d16 and SiT-XL/2, initialising both models with ImageNet-pretrained checkpoints. [^3] VAR: https://github.com/FoundationVision/VAR, SiT: https://github.com/sihyun-yu/REPA (Yu et al., 2025) The base architectures are left unchanged. For VAR we add new learnable positional embeddings for the control tokens and a shared additive “level” embedding for all control tokens, whilst for SiT we only add a new “level” embedding for the control tokens, reusing the original sinusoidal positional embeddings. For SiT, we set diffusion time $t=0$ (clean data) for control tokens. We finetune for 10 epochs with control conditioning with batch size 256 ($\sim$50K iterations) using the original optimisers and hyperparameters. Following the original papers we linearly increase guidance scale $\gamma_{y}$ from zero over generation scales for VAR, whilst keeping it constant for SiT. CFG is performed by only omitting class conditioning and not control conditioning. VAR and SiT use different image crops, leading to small differences in visualisations.
Results.
model
(sampling params)
control
conditioning
#params gen. quality ctrl consistency inf. cost
@@ Results.
VAR-d16
CFG=2.5, temp=1,
top-pp=0.96, top-kk=900
depth prefill 310M 3.30 207 5.95 165 31.3 0.31 12
@@
SiT-XL/2
CFG=1.5, Euler ODE,
steps=64
depth prefill 675M 2.78 218 5.43 170 29.3 3.4 0.35
@@
(a)(b)(c)(d)
Figure 2: Left: Examples of conditional generations using prefilling. Additional examples can be found in Appendix˜B. Right: Results for our prefill baseline where each model is finetuned from an ImageNet-pretrained base model. “*” indicates results copied from Li et al. (2025; 2024b). With default sampling parameters, prefilling is able to (a) generate higher quality images than recent transformer-based approaches with (b) comparable control consistency. We note the (c) relatively low overhead for SiT, as KV-caching means the control tokens use up only a single forward pass. (d) Quantising the control input with VQVAE hurts consistency for VAR, without affecting generation quality.
@@
Takeaway: Prefilling (Fig.˜1) is a simple, general and effective baseline for spatial control with transformers.
We note that our result goes against the results in ControlAR (Li et al., 2025) that suggest that prefilling is a poor choice (they discard this design choice and opt for additive injection of control information, which is what we report in Fig.˜2). We hypothesise that Li et al. (2025)’s choices to 1) vector quantise the control input, destroying information, and 2) enforce a triangular causal attention mask, limiting attention between control tokens, may have hurt performance. We perform our own experiment with VAR’s VQVAE (Fig.˜2) confirming the negative effect of quantising the control input. Although avoiding quantisation may sound obvious, AR models are often trained using token indices as inputs by default (Sun et al., 2024), meaning that “turning off” quantisation may lead to additional implementation complexity.
@@ 5.1 Control Guidance (ctrl-G)
Consider the idea of CFG introduced in Eq.˜6; it can be easily extended to the spatial control condition $\bm{c}$, to concentrate sampling density in regions where $p(\bm{c}|\bm{z})$ is implicitly higher, potentially improving control consistency,
@@
Applying ctrl-G to SiT quickly runs into an issue – generations become highly saturated, shown in Fig.˜3. Thankfully, Sadat et al. (2025) recently propose projected guidance (proj-G) to deal with contrast/saturation issues arising in large CFG scenarios. Applying their approach to ctrl-G, we get,
@@
Takeaway: For flow/diffusion models, control guidance (Eq.˜9) may lead to oversaturation. Applying projected guidance (Sadat et al., 2025) (i.e. only using the orthogonal component Eq.˜11) assuages this issue.
@@ Results.
Fig.˜4 shows how image generation quality and control consistency vary with both CFG ($\gamma_{y}$) and ctrl-G ($\gamma_{c}$). Compared to the default settings used in Fig.˜2, we find that increasing CFG generally improves both IS and FID (although too high a scale may hurt FID), but slightly decreases consistency. Interestingly, in the case of VAR, higher CFG improves FID over the no-control model ($\sim 5.7$ to $\sim 5.1$), demonstrating that providing a weaker model with spatial cues can boost generation quality. On the other hand, ctrl-G is able to drastically improve consistency, but this is traded off against worse generation quality. This is visualised in the left of Fig.˜3, where we see that the canny edge map is better conformed to, but at the cost of over-emphasised edges and incongruous generation for larger ctrl-G. Intuitively, ctrl-G only targets the probability of the control given the generation $p(\bm{c}|\bm{z})$ (Eq.˜8), which may not align with image quality. This trade-off aligns with the results reported for image editing in (Brooks et al., 2023). We note that the F1 values of $\sim$50 for SiT are significantly higher than the values commonly found in the literature (Li et al., 2025; Xiao et al., 2025), demonstrating the large effect that this generation-time adjustment has on consistency. Additionally, ctrl-G incurs the extra inference cost of needing to run the no-control model $p_{\bm{\theta}}(\bm{z})$, reducing latency and throughput: [^6] We use the pretrained ImageNet model for $p_{\bm{\theta}}(\bm{z})$ in series with the control-prefill model. Better latency and parameter efficiency could be achieved by training the control-prefill model with empty control $\bm{c}=\varnothing$ and parallelising like CFG, although the FLOPs increase would still remain.
@@
Finally we note that in ControlVAR (Li et al., 2024b), when performing guidance (with class and control conditioning), all guidance scales $\gamma$ are set to be equal. However, Fig.˜4 shows that this setting may lead to excessive degradation in generation quality, demonstrating the importance of separately tuning $\gamma_{y}$ (CFG) and $\gamma_{c}$ (ctrl-G).
@@ 5.2 Sampling Distribution Truncation
[…]
effectively discarding the tail 1p1-p of the softmax. Intuitively, one would expect the most probable vector codes, predicted by the model’s softmax at each token location, to best conform to the spatial control condition.
@@ Results.
Figs.˜3, 5 and 8 show how top-$p$ and temperature scaling affect generation quality and control consistency alongside CFG. We see that CFG behaves similarly to Fig.˜4, generally improving generation quality and slightly hurting consistency. Notably, truncation via top-$p$ is able to improve both generation quality and control consistency (although excessive truncation seems to hurt FID likely due to reduced diversity). Fig.˜8 (Appendix˜A) further shows that the consistency improvements can be stacked with ctrl-G. This demonstrates that top-$p$ is an important parameter to adjust for softmax-based generative models, providing improvements in all fronts without any additional cost. In contrast, temperature scaling is unable to meaningfully improve consistency, leading to a small increase for canny, and no improvement for depth, whilst simultaneously hurting generation quality. Compared to the results for ctrl-G, this suggests that naively scaling the score (without changing its direction) is insufficient to improve consistency. One possible intuition, is that score temperature scaling occurs uniformly over the spatial dimensions, i.e. it is a global transformation, and so it may have limited spatial/local influence, whilst ctrl-G varies in space. On the other hand, top-$p$ explicitly truncates at each token location, allowing it to influence spatial consistency.
@@
We finally note that Figs.˜4, 5 and 8 demonstrate that meaningfully improved performance can be easily obtained by the sampling adjustments explore in this section (compared to the default settings used in Fig.˜2).
@@ 6.1 Adapter Architectures
Add adapter. Illustrated in the right of Fig.˜6, this approach adapts the ControlNet (Zhang et al., 2023) architecture to transformers, like in Chen et al. (2024a), training a copy of the base model to additively inject spatial conditioning information at each layer, via zero-initialised linear layers, whilst the original weights are performing generation. Inference for each generation token is matched by inference on a corresponding control token.
Attention/prefill adapter. Illustrated in the centre of Fig.˜6, this approach retains the prefill paradigm of conditioning from the baseline, and trains a separate copy of the transformer blocks to be used during condition prefilling, leaving generation to the original model weights. Although the adapter and base model have different attention weight matrices, generation tokens are able to attend back to the keys and values produced by the control adapter.